OpenStudy (anonymous):

Could someone help me on this question please, Simplify: squareroot (18x^8y^3z^2)

OpenStudy (anonymous):

\[\sqrt{18x^{8}y^{3}z^{2}}=3x^{4}yz \sqrt{2y}\]

OpenStudy (anonymous):

Thanks, though how did you get the squareroot of 2y?

OpenStudy (anonymous):

Ohh wait i get it now, thanks a lot for your help!

OpenStudy (anonymous):

Actually, is there a possibility that you could show me some steps? I'm kind of confused, thanks.

OpenStudy (anonymous):

\[\sqrt{18x^8y^3z^2}\]\[\sqrt{18}\times \sqrt{x^8}\times \sqrt{y^3} \times \sqrt{z^2}\]\[\sqrt{2 \times 9}\times x^{8\times \frac{1}{2}} \times y^{3\frac{1}{2}} \times z =\]\[2\sqrt{3}\times x^{4} \times y^{\frac{3}{2}} \times z =\]\[2\sqrt{3 y^3}zx^4\]

OpenStudy (anonymous):

though the answer is 3x^4yz squareroot 2y , but thanks for the help

OpenStudy (anonymous):

Oops, it is\[3\sqrt{2y^3}zx^4\]